# Higher-order-ReLU-KANs (HRKANs) for solving physics-informed neural networks (PINNs) more accurately, robustly and faster

### v1

@@ 4.2 Burgers Equation with viscosity
We solve for the ground-truth solution using Fast Fourier Transform method [46]. As we have not included the implementation of grid extension in ReLU-KAN and HRKAN, to make the comparison between KAN, ReLU-KAN and HRKAN fair, we perform a change of variable $y=\frac{x+5}{4}$ to transform the domain into square domain $[0,2.5]\times[0,2.5]$. The resulting PDE is
@@
We set hyperparameters $\alpha=0.05,N_{\text{pde}}=10000,N_{\text{bc1}}=100,N_{\text{bc2}}=100,N_{\text{ic}}=100$. The train-set data points on the interior are generated randomly while those for boundary and initial conditions are generated uniformly. The structure of the KAN, ReLU-KAN and HRKAN are all of $[2,3,3,3,1]$ with same support for all the bases as $g=7$ and $k=3$. For all three of them, we use the Adam optimizer, set the number of epoch to 3000 and train on the same GPU. We run the same experiment for 10 times.
As said in the previous experiment, we keep the implementation of grid extension in KAN unchanged.
@@
We generate test-set data with a grid with grid size $200\times 200$.
@@
  • From figures 6 and 7 and table 2, it is obvious that the solutions learnt by HRKAN achieve the highest fitting accuracy (test MSE), while the accuracy of the solutions learnt by ReLU-KAN and KAN is relatively lower.
@@
  • The average test MSEs attained by ReLU-KAN and KAN are 0.575% and 1.48% respectively, whereas the average test MSE attained by HRKAN is 0.229%, demonstrating a stronger fitting ability of HRKANs.
  • Also, the standard deviations of test MSEs over the 10 runs are the smallest for HRKAN, at the level of $0.861\%$, where that of ReLU-KAN and KAN are $2.23\%$ and $5.85\%$. This suggests the stronger robustness in the learning of HRKANs.
  • In figure 7, the test MSE converge at the fastest speed for HRKAN, approaching the minimum at around epoch 1500, whereas that of ReLU-KAN and KAN are still on the way of slowly converging at epoch 3000. This suggests the fast convergence speed of HRKANs. The flutuation of the training and test loss of HRKAN after epoch 1250 may imply overfitting of HRKANs after epoch 1250 to the "discontinuous" solution. The relatively more fluctuating training loss and test loss may also due to the lack of regularization tools in ReLU-KANs and HRKANs, compared to KANs.
  • Lastly, we can see that the training time of both HRKAN and ReLU-KAN is much smaller than KAN, exhibiting the efficiency of matrix operations in HRKAN and ReLU-KAN. As mentioned before, there is possibly extra overhead in computation for KANs because of its steps of grid extension and sparsification, but we expect such steps won’t contribute to most of the computation time for KANs, so our judgement remains valid.
@@ 5 Conclusion
In this paper, we proposed a new basis of activation functions for KAN, namely, Higher-order-ReLU. Such Higher-order-ReLU is (1) simpler than the B-Spline basis functions in the original KAN, (2) allows matrix-based operations for fast GPU computing, (3) possesses non-zero and smooth higher-order derviatives essential to Physics-informed neural networks (PINNs). We evaluated Higher-order-ReLU KANs (HRKANs) on two typical PDEs, namely, a linear Poisson equation, and a non-linear Burgers’ equation with viscosity. The detailed experiments exhibit the higher fitting accuracy, stronger robustness and faster convergence speed. Not only can HRKANs find solutions of PDEs, we expect it to have strong potential in further extension like (1) identifying coefficients of PDEs, (2) finding the optimal control of PDEs, and (3) explaining a neural operator etc.

### v2

@@ 4.2 Burgers’ Equation with viscosity
We solve for the ground-truth solution using Fast Fourier Transform method [46]. As we have not included the implementation of grid extension in ReLU-KANs and HRKANs, to make the comparison between KANs, ReLU-KANs and HRKANs fair, we perform a change of variable $y=\frac{x+5}{4}$ to transform the domain into a square domain $[0,2.5]\times[0,2.5]$. The resulting PDE is
@@
In the train-set, the number of interior points is $N_{\text{pde}}=10000$, and the number of boundary points on the 2 boundaries and the number of points for the initial conditions are $N_{\text{bc1}}=100,N_{\text{bc2}}=100$ and $N_{\text{ic}}=100$ respectively. The train-set data points on the interior are generated randomly with uniform distribution while those for the boundary and initial conditions are generated uniformly. We generate test-set data with a grid with grid size $200\times 200$.
@@
We set hyperparameter $\alpha=0.05$.
The structure of the KAN, ReLU-KAN and HRKAN are all of $[2,3,3,3,1]$ with same support for all the bases as $g=7$ and $k=3$. For all three of them, we use the Adam optimizer, set the number of epoch to 3000 and train on the same GPU. We run the same experiment for 10 times. Same as in the previous experiment, we keep the implementation of grid extension and sparsification in KANs unchanged.
@@
  • From figures 6 and 7 and table 2, it is obvious that the solutions learnt by HRKANs achieve the highest fitting accuracy (test MSE), while the accuracies of the solutions learnt by ReLU-KANs and KANs are relatively lower.
@@
  • The mean test MSEs attained by ReLU-KANs and KANs are 0.575% and 1.48% respectively, whereas the mean test MSE attained by HRKANs is 0.229%, demonstrating a stronger fitting ability of HRKANs.
  • Also, the standard deviations of test MSEs over the 10 runs are the smallest for HRKAN, at the level of $0.861\%$, whereas those of ReLU-KAN and KAN are $2.23\%$ and $5.85\%$. This suggests the stronger robustness in the learning of HRKANs.
  • In figure 7, the test MSE converges at the fastest speed for HRKAN, approaching the minimum at around epoch 1500, whereas those of ReLU-KAN and KAN are still on the way of slowly converging at the end of the training process. This suggests the fast convergence of HRKANs. The fluctuation of the training and test loss of HRKANs after epoch 1250 may imply over-fitting of HRKANs after epoch 1250 to the "discontinuous" solution. The relatively more fluctuating training loss and test loss in ReLU-KANs and HRKANs may also due to the lack of regularization tools in these two models, compared to KANs.
  • Lastly, we can see that the training time of both HRKAsN and ReLU-KANs is much smaller than KAN, exhibiting the efficiency of matrix operations in HRKANs and ReLU-KANs. As mentioned before, there is possibly extra overhead in computation for KANs because of its steps of grid extension and sparsification, but we expect such steps won’t contribute to most of the computation time for KANs, so our judgement remains valid.
@@ 5 Conclusion
In this paper, we proposed a new basis of activation functions for KAN, namely, Higher-order-ReLU (HR). Such Higher-order-ReLU is (1) simpler than the B-Spline basis functions in the original KANs, (2) allows matrix-based operations for fast GPU computing, (3) possesses smooth and non-zero higher-order derivatives essential to Physics-informed neural networks (PINNs). We evaluated Higher-order-ReLU-KANs (HRKANs) on two representative and famous PDEs, namely, a linear Poisson equation, and a non-linear Burgers’ equation with viscosity. The detailed experiments exhibit the higher fitting accuracy, stronger robustness and faster convergence of HRKANs. Not only can HRKANs find solutions of PDEs, we expect HRKANs to have strong potential in directions like (1) identifying coefficients of PDEs, (2) finding the optimal control of PDEs, and (3) explaining a neural operator etc.